# Avoid Catastrophic Forgetting with Rank-1 Fisher from Diffusion Models

### v1

@@ Abstract
Catastrophic forgetting remains a central obstacle for continual learning in neural models. Popular approaches—replay and elastic weight consolidation (EWC)—have limitations: replay requires a strong generator and is prone to distributional drift, while EWC implicitly assumes a shared optimum across tasks and typically uses a diagonal Fisher approximation. In this work, we study the gradient geometry of diffusion models, which can already produce high-quality replay data. We provide theoretical and empirical evidence that, in the low signal-to-noise ratio (SNR) regime, per-sample gradients become strongly collinear, yielding an empirical Fisher that is effectively rank-1 and aligned with the mean gradient. Leveraging this structure, we propose a rank-1 variant of EWC that is as cheap as the diagonal approximation yet captures the dominant curvature direction. We pair this penalty with a replay-based approach to encourage parameter sharing across tasks while mitigating drift. On class-incremental image generation datasets (MNIST, FashionMNIST, CIFAR-10, ImageNet-1k), our method consistently improves average FID and reduces forgetting relative to replay-only and diagonal-EWC baselines. In particular, forgetting is nearly eliminated on MNIST and FashionMNIST and is roughly halved on ImageNet-1k. These results suggest that diffusion models admit an approximately rank-1 Fisher. With a better Fisher estimate, EWC becomes a strong complement to replay: replay encourages parameter sharing across tasks, while EWC effectively constrains replay-induced drift.
@@ 1 Introduction
The task of continual learning aims to train neural models on a stream of tasks without revisiting the full past data. A long–standing obstacle is catastrophic forgetting—when learning new tasks drastically degrades performance on earlier ones (McCloskey & Cohen, 1989; French, 1999; Parisi et al., 2019). In contrast, humans exhibit a striking robustness to interference, partly due to systems-level mechanisms such as memory consolidation and replay in the hippocampal–neocortical loop (McClelland et al., 1995; McGaugh, 2000; Foster & Wilson, 2006). These observations have inspired a family of continual learning methods that explicitly encode consolidation or replay in training deep networks on continuous tasks.
Two representative approaches are elastic weight consolidation (EWC) (Kirkpatrick et al., 2017) and generative replay (Shin et al., 2017). EWC constrains parameter changes with a quadratic penalty weighted by an estimate of the Fisher information at a previously learned task, constraining updates to remain near parameter directions that support old tasks (Kirkpatrick et al., 2017; Schwarz et al., 2018). Intuitively, EWC behaves like an online consolidation process that selectively “stiffens” important parameters. Generative replay maintains a generator and distills samples from past tasks while learning new ones, thereby approximating rehearsal without storing original data (Shin et al., 2017; van de Ven & Tolias, 2018). However, replay inherits the generator’s imperfections and can amplify distributional shift. EWC in practice relies on a diagonal Fisher approximation, which neglects cross-parameter correlations and struggles to find a shared parameter space between tasks in overparameterized models, particularly when tasks have disjoint optimum. These observations suggest a complementary pairing: replay can encourage parameter sharing by exposing shared data support, while a stronger approximation to Fisher information can constrain updates and mitigate replay’s residual shift.
In this work, we study the gradient behavior of diffusion models (Ho et al., 2020b; Nichol & Dhariwal, 2021), which are already capable of generating high quality replay data. Our starting point is the observation that diffusion models admit a tractable gradient structure when the signal-to-noise ratio (SNR) is lower (at later timesteps). As a model converges, the per-sample gradients $g$ become approximately collinear with their mean $u$. This makes the empirical Fisher $F$ effectively rank-1:
@@
Our main contributions are: (1) We provide both theoretical and empirical characterizations of Fisher information geometry in diffusion models, showing that low SNR induces a near rank-1 Fisher aligned with the mean gradient. (2) We propose a practical rank-1 EWC penalty that is as cheap as a diagonal penalty but captures more curvature information for diffusion models. (3) We demonstrate that combining rank-1 EWC with distillation-based replay substantially reduces forgetting in continual image generation tasks, improving generation fidelity and stability across long horizons. On MNIST, FashionMNIST, CIFAR-10, and ImageNet-1k, our method consistently outperforms replay-only and diagonal Fisher baselines. In particular, forgetting is nearly eliminated on MNIST and FashionMNIST and is roughly halved on ImageNet-1k, the longest-horizon setting, relative to baselines.
@@ 2.1 Diffusion models
Diffusion models are a family of generative models that define a fixed forward noising process and learn a reverse (denoising) process that maps Gaussian noise to data (Sohl-Dickstein et al., 2015; Ho et al., 2020a). The forward chain corrupts data $x_{0}\sim q_{0}$ via a Markov process
@@
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, $\beta_{t}\in(0,1)$ is the noise level, and $t\in\{0\cdots N\}$ is the discrete forward noising process timestep. The reverse transitions $p_{\theta}(x_{t-1}\!\mid\!x_{t})$ are parameterized by a neural network $\varepsilon_{\theta}$ and trained by maximizing a variational lower bound (ELBO). In practice, the ELBO reduces to a reweighted denoising loss (Ho et al., 2020a):
@@
This surrogate loss can be interpreted as a score-based generative modeling that estimates the score $\nabla_{x_{t}}\log q_{t}(x_{t})$ at each timestep, where $q_{t}(x_{t})=\int q(x_{t}\mid x_{0})q_{0}(x_{0})dx_{0}$, and then integrates a reverse-time SDE/ODE to sample (Song & Ermon, 2019; Song et al., 2021b; Vincent, 2011; Hyvärinen, 2005): $\nabla_{x_{t}}\log q_{t}(x_{t})=-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\,\mathbb{E}[\varepsilon\mid x_{t}],$ and the model’s score estimate follows from $s_{\theta}(x_{t},t)\;=\;-\tfrac{1}{\sqrt{1-\bar{\alpha}_{t}}}\,\varepsilon_{\theta}(x_{t},t)$, where $\varepsilon_{\theta}(x_{t},t)=\mathbb{E}[\varepsilon\mid x_{t}]$ at model optimum.
@@ 2.2 Elastic weight consolidation
Elastic weight consolidation mitigates catastrophic forgetting by regularizing parameter updates using information about how important each weight was to previously learned tasks (Kirkpatrick et al., 2017). It casts continual learning as approximate Bayesian updating: for tasks 1:T1{:}T, the posterior factors as $p(\theta\mid\mathcal{D}_{1:T})\propto p(\mathcal{D}_{T}\mid\theta)\,p(\theta\mid\mathcal{D}_{1:T-1})$. EWC approximates the previous-task posterior $p(\theta\mid\mathcal{D}_{1:T-1})$ with a Laplace approximation around the prior optimum $\theta^{\star}_{T-1}$, yielding $\log p(\theta\mid\mathcal{D}_{1:T-1})~\approx~\frac{1}{2}\,(\theta-\theta^{\star}_{T-1})^{\top}F^{(T-1)}(\theta-\theta^{\star}_{T-1}),$ where $F^{(T-1)}$ is the Fisher information evaluated at $\theta^{\star}_{T-1}$. Plugging this quadratic surrogate into the negative log-posterior for task $k$ gives the EWC objective
@@ 2.3 Continual learning with generative models
Continual learning trains a model on a sequence of tasks, posing the dual challenge of adapting to distribution shift while preserving knowledge from earlier tasks. Broadly, continual learning approaches fall into (i) regularization-based constraints on parameter drift (e.g., EWC) and (ii) replay-based strategies that rehearse past knowledge (Kirkpatrick et al., 2017; van de Ven & Tolias, 2018). For generative models, especially diffusion models, replay is particularly natural because the model can synthesize high-fidelity samples for rehearsal (Ho et al., 2020a).
DDGR uses a diffusion generator with class conditioning to synthesize exemplars of prior tasks, and diffusion-based replay has been adapted to dense prediction (segmentation, detection) using task-specific guidance or pseudo-labels (Gao et al., 2023; Chen et al., 2023; Kim et al., 2024). However, naïve replay with continually updated diffusion models can degrade denoising because the reverse process itself drifts across tasks. Recent work on generative distillation mitigates this by distilling the entire reverse chain ($\theta^{*}_{t-1}\to\theta_{t}$) across timesteps, aligning noise predictions so the continually trained model retains both sample quality and coverage of past tasks (Masip et al., 2025). In this work, we employ generative distillation along with our approach to EWC for a more effective learning.
@@ 3.1 Per-sample gradients align with their mean
If we let $\varepsilon_{\theta}(x_{t},t)$ be the noise prediction network, the per-sample surrogate loss (Ho et al., 2020a) is:
@@
From a denoise score-matching perspective, let model score $s_{\theta}(x_{t},t):=\nabla_{x_{t}}\log p_{\theta}(x_{t})$ and the true marginal score $s_{t}^{\star}(x_{t}):=\nabla_{x_{t}}\log q_{t}(x_{t})$, where $q_{t}$ denotes the noisy data distribution at time $t$ and let $p_{\theta}$ be the model with parameters $\theta$. The connection to per-sample $\mathcal{L}_{\mathrm{simple}}$ can expressed as follows (Ho et al., 2020a; Vincent, 2011; Song et al., 2021b):
@@
$s_{\theta}(x_{t},t)$ approximates a linear function $s_{\theta}(x_{t},t)\approx x_{t}\theta$ when the model learns to perform scaled identity mapping such that $s_{\theta}(x_{t},t)=x_{t}\gamma_{t}I\approx x_{t}\theta,\>\gamma_{t}\in\mathbb{R}$.
@@
Proposition˜1 suggests that when the SNR is low, $s_{t}^{\star}(x_{t})\approx x_{t}\gamma_{t}$, where $\gamma_{t}=-1/(1-\bar{\alpha_{t}})$, which is a scalar independent of input $x_{t}$. Near convergence, $s_{\theta}(x_{t},t)\approx s_{t}^{\star}(x_{t})\approx x_{t}\gamma_{t}$. Assumption˜1 hypothesizes that $s_{\theta}(x_{t},t)$ could be linear and takes the form of $s_{\theta}(x_{t},t)=x_{t}\theta$. This is plausible in practice because a trivial solution for UNet (Ronneberger et al., 2015) is to directly route the inputs to the output due to the skip connections. At a given $t$ such that the SNR is low, $s_{\theta}(x_{t},t)\approx x_{t}\theta\approx x_{t}\gamma_{t}I$. As a result, the gradient of $s_{\theta}(x_{t},t)$ can be understood as the direction that moves $\theta$ to $\gamma_{t}I$. Since $s_{\theta}(x_{t},t)\approx\theta x_{t}$ is linear, the directional change in $\theta$ at each $x_{t}$ is collinear to $\theta-\gamma_{t}I$ independent of $x_{t}$, so the per-sample gradients are collinear with each other, and hence collinear to their mean. We leave the complete proof in Section˜B.2.
@@ 3.2 Empirical validation
Theorem˜1 suggests that, under model convergence, diffusion models approximate a rank-1 Fisher information matrix for their gradients as the SNR decreases (i.e., as the forward process timestep increases). To empirically verify this result, we train a small diffusion model on MNIST (Lecun et al., 1998), whose full Fisher matrix fits in GPU memory, and analyze its gradient behavior across timesteps $t=100,200,\dots,900$. For each $t$, we sample $1024$ data points and compute their gradients. Experiment and model details can be found in Appendix C.
@@ 3.3 EWC penalty with rank-1 Fisher
With the rank-1 approximation using Equation 5, we derive a practical EWC penalty term without forming a full Fisher matrix at each timestep and only use the gradients from the model. For simplicity of notation, let $g=g(\theta;x_{t})$ and $\mu=\mu(\theta)=\mathbb{E}[g]$. From the empirical results in Figure 3(a), the expectation averages both over data and timesteps. With $F(\theta)=\mathbb{E}\big[gg^{\top}\big],$ we have

### v2

@@ Abstract
Catastrophic forgetting remains a central obstacle for continual learning in neural models. Popular approaches—replay and elastic weight consolidation (EWC)—have limitations: replay requires a strong generator and is prone to distributional drift, while EWC implicitly assumes a shared optimum across tasks and typically uses a diagonal Fisher approximation. In this work, we study the gradient geometry of diffusion models, which can already produce high-quality replay data. We provide theoretical and empirical evidence that, in the low signal-to-noise ratio (SNR) regime, per-sample gradients become strongly collinear, yielding an empirical Fisher that is effectively rank-1 and aligned with the mean gradient. Leveraging this structure, we propose a rank-1 variant of EWC that is as cheap as the diagonal approximation yet captures the dominant curvature direction. We pair this penalty with a replay-based approach to encourage parameter sharing across tasks while mitigating drift. On class-incremental image generation datasets (MNIST, FashionMNIST, CIFAR-10, ImageNet-1k), our method consistently improves average FID and reduces forgetting relative to replay-only and diagonal-EWC baselines. In particular, forgetting is nearly eliminated on MNIST and FashionMNIST and is more than halved on ImageNet-1k. These results suggest that diffusion models admit an approximately rank-1 Fisher. With a better Fisher estimate, EWC becomes a strong complement to replay: replay encourages parameter sharing across tasks, while EWC effectively constrains replay-induced drift.
@@ 1 Introduction
The task of continual learning aims to train neural models on a stream of tasks without revisiting the full past data. A long–standing obstacle is catastrophic forgetting—when learning new tasks drastically degrades performance on earlier ones (McCloskey and Cohen, 1989; French, 1999; Parisi and others, 2019). In contrast, humans exhibit a striking robustness to interference, partly due to systems-level mechanisms such as memory consolidation and replay in the hippocampal–neocortical loop (McClelland et al., 1995; McGaugh, 2000; Foster and Wilson, 2006). These observations have inspired a family of continual learning methods that explicitly encode consolidation or replay in training deep networks on continuous tasks.
Two representative approaches are elastic weight consolidation (EWC) (Kirkpatrick et al., 2017) and generative replay (Shin et al., 2017). EWC constrains parameter changes with a quadratic penalty weighted by an estimate of the Fisher information at a previously learned task, constraining updates to remain near parameter directions that support old tasks (Kirkpatrick et al., 2017; Schwarz and others, 2018). Intuitively, EWC behaves like an online consolidation process that selectively “stiffens” important parameters. Generative replay maintains a generator and distills samples from past tasks while learning new ones, thereby approximating rehearsal without storing original data (Shin et al., 2017; van de Ven and Tolias, 2018). However, replay inherits the generator’s imperfections and can amplify distributional shift. EWC in practice relies on a diagonal Fisher approximation, which neglects cross-parameter correlations and struggles to find a shared parameter space between tasks in overparameterized models, particularly when tasks have disjoint optimum. These observations suggest a complementary pairing: replay can encourage parameter sharing by exposing shared data support, while a stronger approximation to Fisher information can constrain updates and mitigate replay’s residual shift.
In this work, we study the gradient behavior of diffusion models (Ho et al., 2020; Nichol and Dhariwal, 2021), which are already capable of generating high quality replay data. Our starting point is the observation that diffusion models admit a tractable gradient structure when the signal-to-noise ratio (SNR) is lower (at later timesteps). As a model converges, the per-sample gradients $g$ become approximately collinear with their mean $u$. This makes the empirical Fisher $F$ effectively rank-1:
@@
Our main contributions are: (1) We provide both theoretical and empirical characterizations of Fisher information geometry in diffusion models, showing that low SNR induces a near rank-1 Fisher aligned with the mean gradient. (2) We propose a practical rank-1 EWC penalty that is as cheap as a diagonal penalty but captures more curvature information for diffusion models. (3) We demonstrate that combining rank-1 EWC with distillation-based replay substantially reduces forgetting in continual image generation tasks, improving generation fidelity and stability across long horizons. On MNIST, FashionMNIST, CIFAR-10, and ImageNet-1k, our method consistently outperforms replay-only and diagonal Fisher baselines. In particular, forgetting is nearly eliminated on MNIST and FashionMNIST and is more than halved on ImageNet-1k, the longest-horizon setting, relative to baselines.
@@ 2.1 Diffusion models
Diffusion models are a family of generative models that define a fixed forward noising process and learn a reverse (denoising) process that maps Gaussian noise to data (Sohl-Dickstein et al., 2015; Ho et al., 2020). The forward chain corrupts data $x_{0}\sim q_{0}$ via a Markov process
@@
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, $\beta_{t}\in(0,1)$ is the noise level, and $t\in\{0\cdots N\}$ is the discrete forward noising process timestep. The reverse transitions $p_{\theta}(x_{t-1}\!\mid\!x_{t})$ are parameterized by a neural network $\varepsilon_{\theta}$ and trained by maximizing a variational lower bound (ELBO). In practice, the ELBO reduces to a reweighted denoising loss (Ho et al., 2020):
@@
This surrogate loss can be interpreted as a score-based generative modeling that estimates the score $\nabla_{x_{t}}\log q_{t}(x_{t})$ at each timestep, where $q_{t}(x_{t})=\int q(x_{t}\mid x_{0})q_{0}(x_{0})dx_{0}$, and then integrates a reverse-time SDE/ODE to sample (Song and Ermon, 2019; Song et al., 2021b; Vincent, 2011; Hyvärinen, 2005): $\nabla_{x_{t}}\log q_{t}(x_{t})=-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\,\mathbb{E}[\varepsilon\mid x_{t}],$ and the model’s score estimate follows from $s_{\theta}(x_{t},t)\;=\;-\tfrac{1}{\sqrt{1-\bar{\alpha}_{t}}}\,\varepsilon_{\theta}(x_{t},t)$, where $\varepsilon_{\theta}(x_{t},t)=\mathbb{E}[\varepsilon\mid x_{t}]$ at model optimum.
@@ 2.2 Elastic weight consolidation
Elastic weight consolidation (Kirkpatrick et al., 2017) mitigates catastrophic forgetting by regularizing parameter updates using information about how important each weight was to previously learned tasks. It casts continual learning as approximate Bayesian updating: for tasks 1:T1{:}T, the posterior factors as $p(\theta\mid\mathcal{D}_{1:T})\propto p(\mathcal{D}_{T}\mid\theta)\,p(\theta\mid\mathcal{D}_{1:T-1})$. EWC approximates the previous-task posterior $p(\theta\mid\mathcal{D}_{1:T-1})$ with a Laplace approximation around the prior optimum $\theta^{\star}_{T-1}$, yielding $\log p(\theta\mid\mathcal{D}_{1:T-1})~\approx~\frac{1}{2}\,(\theta-\theta^{\star}_{T-1})^{\top}F^{(T-1)}(\theta-\theta^{\star}_{T-1}),$ where $F^{(T-1)}$ is the Fisher information evaluated at $\theta^{\star}_{T-1}$. Plugging this quadratic surrogate into the negative log-posterior for task $k$ gives the EWC objective
@@ 2.3 Continual learning with generative models
Continual learning trains a model on a sequence of tasks, posing the dual challenge of adapting to distribution shift while preserving knowledge from earlier tasks. Broadly, continual learning approaches fall into (i) regularization-based constraints on parameter drift (e.g., EWC) and (ii) replay-based strategies that rehearse past knowledge (Kirkpatrick et al., 2017; van de Ven and Tolias, 2018). For generative models, especially diffusion models, replay is particularly natural because the model can synthesize high-fidelity samples for rehearsal (Ho et al., 2020).
DDGR uses a diffusion generator with class conditioning to synthesize exemplars of prior tasks, and diffusion-based replay has been adapted to dense prediction (segmentation, detection) using task-specific guidance or pseudo-labels (Gao and others, 2023; Chen et al., 2023; Kim et al., 2024). However, naïve replay with continually updated diffusion models can degrade denoising because the reverse process itself drifts across tasks. Recent work on generative distillation mitigates this by distilling the entire reverse chain ($\theta^{*}_{t-1}\to\theta_{t}$) across timesteps, aligning noise predictions so the continually trained model retains both sample quality and coverage of past tasks (Masip and others, 2025). In this work, we employ generative distillation along with our approach to EWC for a more effective learning.
@@ 3.1 Per-sample gradients align with their mean
If we let $\varepsilon_{\theta}(x_{t},t)$ be the noise prediction network, the per-sample surrogate loss (Ho et al., 2020) is:
@@
From a denoise score-matching perspective, let model score $s_{\theta}(x_{t},t):=\nabla_{x_{t}}\log p_{\theta}(x_{t})$ and the true marginal score $s_{t}^{\star}(x_{t}):=\nabla_{x_{t}}\log q_{t}(x_{t})$, where $q_{t}$ denotes the noisy data distribution at time $t$ and let $p_{\theta}$ be the model with parameters $\theta$. The connection to per-sample $\mathcal{L}_{\mathrm{simple}}$ can expressed as follows (Ho et al., 2020; Vincent, 2011; Song et al., 2021b):
@@
$s_{\theta}(x_{t},t)$ approximates a linear function $s_{\theta}(x_{t},t)\approx A_{\theta}x_{t}$, where $A_{\theta}$ is a linear operator, when the model learns to perform scaled identity mapping such that $s_{\theta}(x_{t},t)=x_{t}\gamma_{t},\>\gamma_{t}\in\mathbb{R}$.
@@
Proposition˜1 suggests that when the SNR is low, $s_{t}^{\star}(x_{t})\approx x_{t}\gamma_{t}$, where $\gamma_{t}=-1/(1-\bar{\alpha_{t}})$, which is a scalar independent of input $x_{t}$. Near convergence, $s_{\theta}(x_{t},t)\approx s_{t}^{\star}(x_{t})\approx x_{t}\gamma_{t}$. Assumption˜1 hypothesizes that $s_{\theta}(x_{t},t)$ could be linear and takes the form of $s_{\theta}(x_{t},t)=A_{\theta}x_{t}$. This is plausible for an autoencoder with nonlinearities optimized with the MSE objective, such as the $\mathcal{L}_{\mathrm{simple}}$ in Ho et al. (2020), as these autoencoders are still “likely” to lie in a PCA linear subspace (Vincent et al., 2010). We leave additional discussion and analysis on alternative autoencoding objectives in Appendix C. At a given $t$ such that the SNR is low, $s_{\theta}(x_{t},t)\approx A_{\theta}x_{t}\approx x_{t}\gamma_{t}$. As a result, the gradient of $s_{\theta}(x_{t},t)$ can be understood as the direction that moves $A_{\theta}$ to $\gamma_{t}I$. Since $s_{\theta}(x_{t},t)\approx A_{\theta}x_{t}$ is linear, the directional change in $A_{\theta}$ at each $x_{t}$ is collinear to $A_{\theta}-\gamma_{t}I$ independent of $x_{t}$, so the per-sample gradients are collinear with each other, and hence collinear to their mean. We leave the complete proof in Section˜B.2.
@@ 3.2 Empirical validation
Theorem˜1 suggests that, under model convergence, diffusion models approximate a rank-1 Fisher information matrix for their gradients as the SNR decreases (i.e., as the forward process timestep increases). To empirically verify this result, we train a small diffusion model on MNIST (Lecun et al., 1998), whose full Fisher matrix fits in GPU memory, and analyze its gradient behavior across timesteps $t=100,200,\dots,900$. For each $t$, we sample $1024$ data points and compute their gradients. Experiment and model details can be found in Appendix D.
@@ 3.3 EWC penalty with rank-1 Fisher
With the rank-1 approximation using Equation 5, we derive a practical EWC penalty term without forming a full Fisher matrix at each timestep and only use the gradients from the model. For simplicity of notation, let $g=g(\theta;x_{t})$ and $\mu=\mu(\theta)=\mathbb{E}[g]$. In implementation, this expectation is taken over the joint sampling process used during diffusion training: we draw a datapoint $x_{0}\sim p_{\mathrm{data}}$, sample a timestep $t\sim p(t)$ (uniform in our experiments unless stated otherwise), construct the noised input $x_{t}$ from $(x_{0},t)$, and then compute $g(\theta;x_{t},t)=\nabla_{\theta}\ell(\theta;x_{t},t)$. Thus, $\mu$ should be understood as $\mu(\theta)=\mathbb{E}_{x_{0},t}\!\left[g(\theta;x_{t},t)\right]$
This comes from the observation that the SNR decreases sharply at early timesteps via Tweedie’s identity, and therefore remains low for the majority of diffusion steps (see Figure 1). Consistent with this, the empirical Fisher is strongly dominated by its leading eigen-direction across timesteps. For instance, we observe $\lambda_{2}/\lambda_{1}<0.4$ already at $t=100$ in Figure 3(c), and the mean gradient at different timesteps is highly collinear with one another in Figure 3(a). This indicates that the rank-1 approximation captures a large portion of the curvature even outside the latest timesteps. Consequently, averaging gradients over timesteps yields a practical surrogate that dilutes potentially residual higher-rank structure from the earliest steps and is dominated by the nearly rank-1 gradients from the remaining timesteps. With $F(\theta)=\mathbb{E}\big[gg^{\top}\big],$ we have